OpenStudy (anonymous):

Please Help Find the slope-intercept equation for the tangent line to the curve: y=ln((1+x)/(1-x)) at x=0.1

OpenStudy (anonymous):

y-1=2x

OpenStudy (anonymous):

how did you do that?

OpenStudy (anonymous):

c 2 points are given and the last thing u need is m (slope) which in this case will be dy/dx which is = 2 in this case then eq of line is y-y1=m(x-x1) y-1=2(x-0) y-1=2x if u understood plz do tell..